OpenStudy (anonymous):

What does parametrization of a curve in vector calculus define? What does parametrization of a curve in vector calculus define? @Mathematics

OpenStudy (jamesj):

It's a way of defining a curve in such a way that you can 'move' along the curve as a function of one variable, that you can think of as time. The best way to think about this is through examples. Here's a parameterization of a circle, where you continue to go round and round the circle: r(t) = (cos(t),sin(t)) Clearly this is a circle because with x = x(t) = cos(t) and y = y(t) = sin(t) x^2 + y^2 = 1 Here's a parameterization of the circle where you go around 5 times as fast: r(t) = (cos(5t),sin(5t)) And here's a parameterization of a spiral in the three dimensional space: r(t) = (cos(t), sin(t), t)